RRRG report for The Peak Express October 2011, Issue 29
Work on Renown continues to be centred on identifying and labelling cables, and establishing exactly what has been cut. Areas looked at so far include, the eight terminal bars in the Thin Man's passage, the Fire Alarm Relay No1, the cubicle base, No1 cab, the water header tank level switch, the Battery Isolation Switch and AWS Relay Box. Whilst the opportunity presented itself the water header tank level switch operation was tested and found to be good and therefore the whole assembly was reinstalled. The auxiliary contacts off the 'proper' class 50 reverser (to be used in 50030) have been temporarily removed, and are being used to provide a pattern for the fabrication of a new mechanism for the second reverser for Renown. This might not sound like a lot of progress, but as I said last month, the job of tracing and labelling all of the wires is a huge one.
In the generator room of both RRRG locomotives the potentiometer stand had been bent at sometime, most likely when the ETH/Auxiliary generator sets were removed. With the aid of some heavy wood and a heavy duty jack this steel frame has been straightened, welded, cleaned, refitted and painted in both locomotives. Also in 50030's generator room the damaged floor mounted conduits have all been removed and the floor cleaned ready for painting. The conduits will be replaced with a mixture of those salvaged from 500023, 50030 and 50040 prior to its scrapping.
Work is also continuing a pace on the complicated task of rebuilding 50030's electrical cubicle. The auxiliary contacts on BMR1, GFR & OIR relays and the Z59 resistor on the board behind the fuse cabinet have been wired. The three remaining sections of 185mm cable to the field divert contactors, that had proved troublesome have now been fitted. This has enabled the three boards that go in front of said cabling in the Thin Man's Passage to be fitted and wired. These boards mount the ETH overvolt relay HOVR & it's attendant resistor Z24, ETH regulator resistor Z23, the ETH earth parallel resistor Z69, ETH detector relay HDR, ETH earth-fault relay HEFR & it's resistor Z62, the ETH overload relay HOR, ETH timing relay HTR & the ETH overload resistor Z70. Prior to fitting these boards, all the above items were removed from them, stripped, cleaned and refitted. The HOVR, Z23, Z24, Z69, Z62 and HDR have been wired in. These are the last boards for the electrical cubicle. When testing Z70, it was found to be open-circuit and therefore useless, so this is another item to be checked out when Andy is next working on 50008; to measure its resistance, so a substitute can hopefully be found.
Two good condition NR28 relays from our stock have had their contact blocks removed, cleaned and reassembled and have been fitted into the cubicle. These are the last two relays to go into Repulse. Andy is also now in the process of checking each of the eight terminal bars in the Thin Man's passage, to make sure all the wiring that should be there is there. The last three cables to the CU5 radiator fan control unit have been laid in. Work is also progressing well in the cabs with the work to repair and replace the runner timbers below the droplights and the subframes around the quarterlights being virtually complete. Attention has now turned to replacing the interior wood surrounds starting with the quarter-lights & drop-lights in 30's No.2 cab. One of the DSD foot pedals has been dismantled, cleaned and freed up as years of no use had stiffened things up a little, the switch assembly surprisingly looked very clean and on testing after a quick clean of the switch contacts worked and tested ok.
On the mechanical side attention has now turned to the brake rigging, the spare brake rigging has been sorted out, although it has been in store for a long time, a lot of this rigging has proved to be in very good condition, work is now progressing on cleaning and un-sticking various parts. The brake rigging on one of the spare bogies has also been removed and assessed, although a lot of this was found to be badly worn, some items have been put to one side for future use. Work is also progressing on the spare bogies to prepare the traction motors for lifting, although the fixing bolts (as with the brake rigging) are proving very tough!!! The plan is for the three traction motors from one bogie and two other spare traction motors to be lifted out and sent to Bowers at the time that we take delivery of the overhauled exhausters and motor blowers. Once overhauled these three motors will be swapped for the three motors on the other spare bogie.
One of the main tasks to have been undertaken on the power unit during the last few months is the torquing up of all the big-end stretch bolts! All 64 of them, this also entailed making a couple of special tools for the job as the power unit internal structural webbing gets in the way of using an ordinary spanner properly. Attention was then turned to the camshafts, which have all been freed up and greased. Work is now progressing on reassembling the valve gear.
Finally our sales stand has had an excellent three weekends out at the Nene Valley Gala, the Seven Valley Gala and the East Lancs Deltic weekend, the only downside to this is that it has had a serious impact on our stock of secondhand items, particularly diesel related books which are down to half a box, and model railway items which we are down to two diesel locos! As such we would be very grateful to anyone that is able to preferably donate, or sell for a reasonable price, any items to the group.